OpenStudy (anonymous):

The lengths of the sides of a triangle are 8, 11, and 14. Find the perimeter of a similar triangle whose longest side measures 21.

OpenStudy (nikvist):

\[\frac{O_2}{O_1}=\frac{a_2}{a_1}=\frac{b_2}{b_1}=\frac{c_2}{c_1}=\frac{21}{14}=1.5\] \[O_2=1.5O_1=1.5(8+11+14)=49.5\]

OpenStudy (anonymous):

can u explain to me how you got that ?

OpenStudy (nikvist):

These triangls are similar. Similarity ratio is 21/14=1.5. Their perimeters are similar also.

OpenStudy (anonymous):

where did u get 1.5 ?

OpenStudy (anonymous):

where did u get that equation ?

OpenStudy (nikvist):

longest side of 2nd triangle : longest side of 1st triangle = 21:14 =1.5

OpenStudy (anonymous):

oh i understand now. thanks